### v2

# Hermitian-symplectic and Kähler structures on degenerate twistor deformations

Abstract Let $(M,\Omega)$ be a holomorphically symplectic manifold equipped with a holomorphic Lagrangian fibration $\pi:M\to B$, and $\eta$ a closed $(1,1)$-form on $B$. Then $\Omega+\pi^{*}\eta$ is a holomorphically symplectic form on a complex manifold which is called the degenerate twistor deformation of $M$. We prove that degenerate twistor deformations of compact holomorphically symplectic Kähler manifolds are also Kähler. First, we prove that degenerate twistor deformations are Hermitian symplectic, that is, tamed by a symplectic form; this is shown using positive currents and an argument based on the Hahn–Banach theorem, originally due to Sullivan. Then we apply a version of Huybrechts’s theorem showing that two non-separated points in the Teichmüller space of holomorphically symplectic manifolds correspond to bimeromorphic manifolds if they are Hermitian symplectic.
1 Introduction
1.1 Degenerate twistor deformations and the Tate–Shafarevich twist
Let $(M,I,\Omega)$ be a holomorphically symplectic manifold, that is, an almost complex manifold equipped with a closed, non-degenerate $(2,0)$-form. Clearly, the almost complex structure $I$ is uniquely determined by the differential form $\Omega$. It is not hard to see that the integrability of the almost complex structure $I$ follows from closedness of $\Omega$ ([V3]). This allows one to define the complex structure in terms of $\Omega$, giving a motivation for the notion of C-symplectic form (2.2).
The notion of degenerate twistor deformation, introduced in [V3], and further explored in [BDV] and [SV], is one of the applications of this approach. Given a holomorphic Lagrangian fibration $\pi:\;M{\>\longrightarrow\>}B$ on a holomorphic symplectic manifold $(M,\Omega)$, we can replace $\Omega$ by $\Omega+t\pi^{*}\eta$, where $\eta$ is a closed $(2,0)+(1,1)$-form on $B$, and $t\in{\mathbb{C}}$. The new form $\Omega_{t}:=\Omega+t\pi^{*}\eta$ is C-symplectic, that is, holomorphic symplectic with respect to a complex structure $I_{t}$ (2.2). This defines a holomorphic family of complex manifolds $(M,I_{t}),t\in{\mathbb{C}}$, which is called the degenerate twistor deformation.
It shares many of its properties with twistor deformations (the ones which come from a hyperkähler rotation); indeed, the degenerate twistor deformation can be obtained as a limit of twistor deformations ([V3, Section 1.2]).
A version of this construction is known as the “Tate–Shafarevich twist”. For elliptic surfaces, it was explored at some length in [FM], and for higher dimensions in [Mar] and [AR]. In [AR] it was shown that the Tate–Shafarevich twist and the degenerate twistor deformation coincide if the Lagrangian fibration has no multiple fibers in codimension 1.
In the present paper, we prove that the degenerate twistor deformation is always Kähler (4.3), answering the question which was originally posed by E. Markman.
Theorem 1.1: Let $M$ be a compact holomorphically symplectic manifold of Kähler type, $\pi:\;M{\>\longrightarrow\>}B$ a Lagrangian fibration, and $\tau:\;{\cal M}{\>\longrightarrow\>}{\mathbb{C}}$ the corresponding degenerate twistor deformation. Then all fibers of $\tau$ admit a Kähler metric.
Proof: 4.3.
In [Mar], Markman has studied the Tate–Shafarevich twist of a Lagrangian fibration on a Hilbert scheme of a K3 surface. Markman calls a compact holomorphically symplectic manifold $(M,\Omega)$ “M-special” if the space $\langle[\operatorname{Re}\Omega],[\operatorname{Im}\Omega]\rangle\subset H^{2}(M,{\mathbb{R}})$ does not contain non-zero rational cohomology classes. He has shown that the Tate–Shafarevich twist of a Lagrangian fibration on a Hilbert scheme of a K3 surface is Kähler, if it is not M-special. In [AR], this result was generalized to all Lagrangian fibrations without multiple fibers in codimension 1. In the present paper we give an independent proof that works for arbitrary compact hyperkähler manifolds without the assumption of non-M-specialty.
1.2 Non-Kähler deformations of hyperkähler manifolds
The earliest publication on Kählerness of holomorphically symplectic manifolds was a preprint of Andrey Todorov [T2]. Using a Harvey-Lawson-type Hahn-Banach argument, Todorov tried to show that a simply connected, compact, holomorphically symplectic manifold with $H^{2,0}(M)={\mathbb{C}}$ is Kähler.
The earliest counterexamples to this statement were obtained in 1994 by D. Guan, [Gu1, Gu2, Gu3, Bo]. These are manifolds, obtained from the Kodaira surface (non-Kahler, holomorphically symplectic complex surface) by a procedure reminding of the one which is used to construct the generalized Kummer variety from a 2-dimensional torus. This manifold is non-Kähler because it contains a complex surface isomorphic to a blow-up of Kodaira surface ([Bo, Corollary 4.10]).
Another example of a non-Kähler holomorphically symplectic manifold comes from birational geometry. [^1] We benefited from a Mathoverflow discussion at http://mathoverflow.net/questions/302649/are-there-non-projective-but-algebraic-hyperkahler-varieties and the information provided by the user YangMills. Recall that Mukai flop is a bimeromorphic transform which takes a Lagrangian ${\mathbb{C}}P^{n}$ in a holomorphically symplectic manifold $M$, and replaces it with another one, obtaining another holomorphically symplectic manifold (say, $M^{\prime}$). The second homology groups of these manifolds are isomorphic, however, the homology class of a rational curve (a line contained in the ${\mathbb{C}}P^{n}$) on $M$ is opposite to that on $M^{\prime}$. If we start from a holomorphically symplectic manifold containing two disjoint Lagrangian planes with homologous lines, and do a flop in only one of these, we will obtain a holomorphically symplectic manifold which contains a complex curve (the union of two lines) homologous to zero. This manifold is Fujiki class C (that is, bimeromorphic to Kähler). For explicit examples of non-Kähler Fujiki class C holomorphic symplectic manifolds and a reference, see [GHJ, Example 21.9].
In the sequel, hyperkähler means “compact holomorphically symplectic manifold of Kähler type”. In [Hu1], D. Huybrechts proved that two bimeromorphic hyperkähler manifolds corresponds to non-separable (in the sense of being non-Hausdorff) points of the corresponding Teichmüller space. Conversely, given two non-separable points $I_{1},I_{2}$ in the Teichmüller space, the hyperkähler manifolds $(M,I_{1})$ and $(M,I_{2})$ are bimeromorphic. Curiously enough, this argument uses the Kählerness in an essential way.
If one drops the Kählerness assumption (or a weaker assumption of existence of the Hodge decomposition on $H^{2}(M)$ as in [Per]; see the discussion of this work below), the proof is still unknown, though we may dare to make a conjecture.
Conjecture 1.2:  Let $(M_{t},\Omega_{t})$ be a smooth family of compact holomorphically symplectic manifolds over a disk $\Delta$. Assume that all fibers except the central one are hyperkähler. Then the central fiber $(M_{0},\Omega_{0})$ is of Fujiki class C.
In 4.2, we prove this conjecture when the central fiber $M_{0}$ is Hermitian symplectic. In [Per], A. Perego discussed 1.2 when $M_{0}$ satisfies $b_{2}(M_{0})=h^{2,0}(M_{0})+h^{1,1}(M_{0})+h^{0,2}(M_{0})$.
Let us indicate where Huybrechts’s proof uses the Kählerness. Consider two deformations $M_{t}$ and $M_{t}^{\prime}$ of a hyperkähler manifold, with $t\in\Delta$. Assume that there is an isomorphism $M_{t}\cong M^{\prime}_{t}$ for all $t\neq 0$. The graph of this isomorphism defines a family of subvarieties $\Gamma_{t}$ in $M_{t}\times M^{\prime}_{t},t\neq 0$. Using Bishop’s compactness theorem, Huybrechts proves that the family $\Gamma_{t}$ can be extended to $t=0$, and the corresponding cycle $\Gamma_{0}$ contains a component that is the graph of a bimeromorphic map $M_{0}\dashrightarrow M^{\prime}_{0}$. When $M_{0}$ is non-Kähler, Bishop’s compactness fails and the argument becomes invalid. However, Bishop’s compactness still holds for Hermitian symplectic manifolds (4.1), and this leads to 4.2.
The application of this argument often brings an error, which was observed by Y.-T. Siu [S1, S2] when he was discussing a paper of A. Todorov [T1] “Applications of the Kähler-Einstein-Calabi-Yau metric to moduli of K3 surfaces.” The same problem occurs later in two important and very interesting papers [Per, A].
As explained by Siu, the gaps occur in the application of E. Bishop’s theorem on limits of complex analytic subvarieties. Suppose that we have two families of holomorphically symplectic manifolds over a disk, isomorphic except in the central fiber; as in Huybrechts’s proof, one considers the limit of the graphs $\Gamma_{t}$ of the isomorphisms between the non-central fibers. The idea is to apply Bishop’s theorem to show that this limit is a complex analytic subvariety with the same fundamental class.
To apply Bishop’s theorem, bounding the volume of the limiting cycle, one needs to choose a Hermitian form on the central fiber, such that the volume of $\Gamma_{t}$ with respect to this form stays bounded for all $t$. Even if we construct a sequence of Kähler metrics on the neighboring fibers whose cohomology classes converge to a cohomology class on the central fiber (such as in [Per, Lemma 2.5] and in [A, Lemma 4.6.1]), this does not imply the volume estimates. Indeed, a cohomology class with positive BBF square on a holomorphic symplectic manifold that is a limit of Kähler classes on neighboring fibers needs not be Kähler itself.
1.2 is a special case of a more ambitious conjecture proposed in [LL].
Conjecture 1.3:  Let $M_{t}$ be a smooth family of compact complex manifolds over a disk $\Delta$. Assume that all fibers except the central one are Kähler. Then the central fiber $M_{0}$ is Fujiki class C.
In [LL], this conjecture is proven under two extra assumptions on a central fiber: (a) $M_{0}$ admits a Hermitian form $\omega$ satisfying $dd^{c}\omega=0$ and $\partial\omega\wedge\overline{\partial}\omega=0$, and (b) $h^{0,2}(M_{0})=h^{0,2}(M_{t})$ for $t\neq 0$.
In [Po], this conjecture is proven when $M_{t}$, $t\neq 0$ are projective, and $h^{0,1}(M_{0})=h^{0,1}(M_{t})$ for $t\neq 0$; in this case, the central fiber is Moishezon. Later, this result was amplified by [LRWW], who proved that a limit of Moishezon manifolds is Moishezon.
1.3 Hermitian symplectic geometry and Streets–Tian conjecture
The original impulse to this paper is an observation due to Peternell: any non-projective Moishezon manifold admits a positive, exact $(n-1,n-1)$-current, see [Pe]. We were trying to use this observation to prove that a degenerate twistor deformation of a hyperkähler manifold cannot be non-projective Moishezon, in hope to extend it further to Fujiki class C, and then try to prove 1.2 in this particular situation.
In 3 we accomplish the first task, indeed, showing that a degenerate twistor deformation of a hyperkähler manifold cannot support a positive, exact $(n-1,n-1)$-current.
Let us illustrate this in a simpler situation, showing that a degenerate twistor deformation cannot admit a complex curve $C$ which is homologous to zero.
Let $\pi:\;M{\>\longrightarrow\>}B$ be a Lagrangian fibration on a hyperkähler manifold $(M,\Omega)$. Then $B$ is projective; let $\omega\in H^{2}(B,{\mathbb{R}})$ be its Kähler class. Since $\int_{C}\pi^{*}\omega=0$, $C$ is contained in the fibers of $\pi$. However, the degenerate twistor deformation does not change the complex structures on the fibers, hence $C$ is a complex curve on the original hyperkähler manifold, which is impossible, because a subvariety of a Kähler manifold cannot be homologous to zero.
In 3 we modify this argument showing that it remains valid for any positive, exact $(n-1,n-1)$-current in place of the current of integration of $C$.
A theorem of Sullivan [Su] implies that a complex manifold admits no positive, exact $(n-1,n-1)$-currents if and only if it is Hermitian symplectic, that is, admits a taming symplectic structure (3).
Streets and Tian have conjectured that Hermitian symplectic manifolds are Kähler, see [ST]. This conjecture was treated at great length, and solved for some classes of complex manifolds (for complex surfaces in [LZ, Theorem 1.2], for twistor spaces in [V2]), but in full generality it remains very difficult. In [EFV, FV] the Streets–Tian conjecture was proven for complex nilmanifolds.
The Streets–Tian conjecture makes sense for almost complex manifolds as well; in this context, it is known as the “Donaldson’s tamed to compatible conjecture”. In [Do, Question 2], Donaldson asked whether an almost complex 4-manifold admitting a taming symplectic form must also admit a compatible symplectic form.
Using 3, we reduce the proof of Kählerness of degenerate twistor deformations to the Streets–Tian conjecture, and in Subsection 4.2, we prove a special case of this conjecture, finishing the proof of Kählerness.
Theorem 1.4: Let $\operatorname{\sf Teich}$ be the Teichmüller space of holomorphically symplectic complex structures, and $I\in\operatorname{\sf Teich}$ a holomorphically symplectic complex structure obtained as a limit $I=\lim_{i}I_{i}$, where $I_{i}\in\operatorname{\sf Teich}$ are holomorphically symplectic complex structures of Kähler type. Assume that $(M,I)$ is Hermitian symplectic. Then $(M,I)$ is also of Kähler type.
Proof: By 4.2, $(M,I)$ is of Fujiki class C. By [Ch, Theorem 0.2], any Hermitian symplectic Fujiki C manifold is of Kähler type.
2 Preliminaries
2.1 Hyperkähler manifolds
We start by recalling the basic facts of the theory of hyperkähler manifolds. For a detailed exposition of this theory see [Hu1].
We will denote by $\mathbb{H}$ the algebra of quaternions and by $Sp(n)$ the group of quaternionic-linear transformations of $\mathbb{H}^{n}$. Let $M$ be a compact $C^{\infty}$ manifold. A Riemannian metric $g$ on $M$ is called hyperkähler if the holonomy group of its Levi–Civita connection $\nabla^{g}$ is contained in $Sp(n)$. The elements of $Sp(n)$ commute with a two-dimensional family of complex structures defined via multiplication by imaginary unit quaternions. Hence by the holonomy principle we get a family of $\nabla^{g}$-parallel complex structures on $M$ called the twistor family.
In this paper we will always assume that hyperkähler metrics are of maximal holonomy, i.e. the holonomy group of $\nabla^{g}$ is isomorphic to $Sp(n)$. Given a complex structure $I$ on $M$, we will say that $I$ is of hyperkähler type if there exists a hyperkähler metric of maximal holonomy $g$ on $M$ such that $\nabla^{g}I=0$. If $I$ is of hyperkähler type and $M$ is simply connected, the manifold $(M,I)$ is usually called irreducible holomorphic symplectic (IHS).
Given a hyperkähler metric $g$ on $M$ one can find $\nabla^{g}$-parallel complex structures $I$, $J$ and $K$ such that $K=IJ=-JI$. Denote by $\omega_{I}$, $\omega_{J}$ and $\omega_{K}$ the corresponding Kähler forms. The 2-form $\Omega=\omega_{J}+\sqrt{-1}\omega_{K}$ is a holomorphic symplectic form on $(M,I)$.
Recall the following facts about the cohomology of a hyperkähler manifold $M$ of maximal holonomy (see [Bea] and [Hu1] for the details). The vector space $H^{2}(M,\mathbb{Q})$ carries a non-degenerate quadratic form $q$ called the Beauville–Bogomolov–Fujiki (BBF) form. It has the following property: there exists a positive constant $c_{M}\in\mathbb{Q}$ such that for all $a\in H^{2}(M,\mathbb{Q})$ the Fujiki relation holds:
$\int_{M}a^{2n}=c_{M}q(a)^{n}.$ (2.1)
We will normalize $q$ in such a way that it is primitive and integral on $H^{2}(M,\mathbb{Z})$. The BBF form has signature $(3,b_{2}(M)-3)$.
Consider the collection $\operatorname{\sf Comp}(M)$ of complex structures of hyperkähler type on $M$ and the group $\operatorname{\sf Diff}^{\circ}(X)$ of diffeomorphisms of $M$ isotopic to the identity. The quotient $\operatorname{\sf Teich}(M)=\operatorname{\sf Comp}(M)/\operatorname{\sf Diff}^{\circ}(M)$ carries the structure of a non-Hausdorff complex manifold and is called the Teichmüller space of $M$. The global Torelli theorem for hyperkähler manifolds [V1] describes the Teichmüller space in terms of the cohomology of $M$. Define the period domain:
$\mathcal{D}=\{x\in\mathbb{P}H^{2}(M,\mathbb{C})\enskip|\enskip q(x)=0,\,q(x,\overline{x})>0\}$
If $I$ is a complex structure of hyperkähler type then the space of $I$-holomorphic two-forms on $M$ is spanned by the symplectic form $\Omega$, and the corresponding point $[\Omega]\in\mathcal{D}$ is called the period of $I$. This defines the period map $\operatorname{\sf Per}\colon\operatorname{\sf Teich}(M)\to\mathcal{D}$. The global Torelli theorem [V1] claims that the restriction of $\operatorname{\sf Per}$ to every connected component of $\operatorname{\sf Teich}$ is the Hausdorff reduction map, i.e. after identifying the non-separated points of $\operatorname{\sf Teich}$ it induces an isomorphism of complex manifolds.
2.2 Degenerate twistor deformations
In this subsection, we briefly recall the definition of a degenerate twistor deformation associated with a holomorphic Lagrangian fibration on a hyperkähler manifold, following [BDV] (see also [V3] and [SV]). Holomorphic Lagrangian fibrations are ubiquitous in hyperkähler geometry, as shown by the following foundational theorem.
Theorem 2.1: (D. Matsushita) Let $\pi\colon M\to B$ be a surjective holomorphic map from a compact hyperkähler manifold $M$ of maximal holonomy to a complex variety $B$ with $0<\dim B<\dim M$. Then $\dim B={1\over 2}\dim M$, and the fibers of $\pi$ are holomorphic Lagrangian.
Proof: [Mat].
A map as in the above theorem is called a holomorphic Lagrangian fibration. The base $B$ is conjectured to be ${\mathbb{C}}P^{n}$, and this is the case in all known examples; when $B$ is smooth, it is biholomorphic to ${\mathbb{C}}P^{n}$, as shown in [Hw].
Definition 2.2:  Let $M$ be a real $4n$-dimensional $C^{\infty}$-manifold. A C-symplectic form on $M$ is a smooth section $\Omega$ of the complexified bundle of 2-forms $\Lambda^{2}M\otimes\mathbb{C}$ such that: $d\Omega=0$, $\Omega^{n+1}=0$ and the form $\Omega^{n}\wedge\overline{\Omega}\mathstrut^{n}$ is pointwise non-vanishing.
Let $\Omega$ be a C-symplectic form on $M$. For a point $x\in M$ let $T^{0,1}_{x}M\subset(T_{x}M)\otimes\mathbb{C}$ be the kernel of $\Omega$, i.e. the subspace consisting of complex tangent vectors $v\in(T_{x}M)\otimes\mathbb{C}$ such that the contraction $v\,\lrcorner\,\Omega$ vanishes.
One verifies (see [BDV], [SV]) that the kernel of $\Omega$ has rank $2n$ at all points of $M$, hence $T^{0,1}M\subset TM\otimes\mathbb{C}$ is a subbundle. Defining $T^{1,0}M$ as the complex conjugate of $T^{0,1}M$, one gets the decomposition $TM\otimes\mathbb{C}=T^{1,0}M\oplus T^{0,1}M$. Using the closedness of the form $\Omega$, one checks that the subbundle $T^{1,0}M$ is closed under the Lie bracket. This defines an integrable almost complex structure on $M$, and therefore by the Newlander–Nirenberg theorem any manifold with a C-symplectic form admits an intrinsically defined complex structure.
Theorem 2.3:  Let $M$ be a compact complex manifold with a holomorphic symplectic form $\Omega$, and let $\pi\colon M\to B$ be a holomorphic Lagrangian fibration. Consider a closed two-form $\eta\in\Lambda^{2}B\otimes\mathbb{C}$ such that $\eta^{0,2}=0$. Then the two-form $\Omega_{t}:=\Omega+t\pi^{*}\eta$ is C-symplectic for any $t\in\mathbb{C}$. If we denote by $I_{t}$ the complex structure on $M$ corresponding to $\Omega_{t}$, then for any $t\in\mathbb{C}$ the map $\pi$ is holomorphic with respect to $I_{t}$ and the fixed complex structure on the base $B$.
Proof: [BDV].
The following property of the complex structures $I_{t}$ from the above theorem appears implicitly in [BDV] and [SV]. Let us record it for the later use.
Lemma 2.4:  In the setting of 2.2 denote by $F$ one of the fibres of $\pi$ and let $x\in F$. Let $V=T_{x}F\otimes\mathbb{C}$ be the complexified tangent space of $F$ at the point $x$. Then the subspace $V_{t}^{0,1}=V\cap T^{0,1}_{I_{t}}M$ does not depend on $t\in\mathbb{C}$.
Proof: Consider $v\in V$. Then $v\in V^{0,1}_{t}$ if and only if $v\,\lrcorner\,\Omega_{t}=0$. Note that $V=\ker(d\pi)_{x}$, therefore $v\,\lrcorner\,\pi^{*}\eta=0$ and $v\,\lrcorner\,\Omega_{t}=v\,\lrcorner\,\Omega$. Hence $V^{0,1}_{t}=V^{0,1}_{0}$, so $V^{0,1}_{t}$ is a fixed subspace of $V$ not depending on $t$.
Assume now that $M$ is a compact simply connected hyperkähler manifold of maximal holonomy, and $\pi\colon M\to B$ a holomorphic Lagrangian fibration over a projective base $B$ (the projectivity of $B$ follows from the other assumptions, see [AC, footnote on page 53]). Let $\eta\in\Lambda^{1,1}B$ be the restriction of the Fubini–Study form under some projective embedding of the base, $\Omega_{t}=\Omega+t\pi^{*}\eta$ the C-symplectic form and $I_{t}$ the corresponding complex structure, where $t\in{\mathbb{C}}$. The manifolds $\mathcal{M}_{t}=(M,I_{t})$ are fibres of the the degenerate twistor deformation of $M$, which is a smooth holomorphic family $f\colon\mathcal{M}\to\mathbb{A}^{1}$ over the affine line, see [V3]. This deformation has a number-theoretic interpretation (“the Shafarevich–Tate family”), see [Mar, AR].
The periods of the manifolds $\mathcal{M}_{t}$ form an affine line in the period domain. More precisely, consider the subspace $W\subset H^{2}(M,\mathbb{C})$ spanned by $\Omega$, $\overline{\Omega}$ and $\pi^{*}\eta$. Then the restriction of the BBF form $q|_{W}$ is a degenerate positive-semidefinite quadratic form with one-dimensional kernel spanned by $\pi^{*}\eta$. The corresponding quadric $Q_{W}$ is the union of two projective lines intersecting at the point $[\pi^{*}\eta]$. The complement to that point in $Q_{W}$ is the disjoint union of two affine lines, one of them passing through $[\Omega]$ and the other through $[\overline{\Omega}]$. The first of the two affine lines is formed exactly by the periods of the manifolds $\mathcal{M}_{t}$, $t\in\mathbb{C}$ of the degenerate twistor family. The other affine line corresponds to the complex-conjugate degenerate twistor family which one obtains replacing $\Omega$ by $\overline{\Omega}$.
2.3 Weakly positive and strongly positive forms
When we speak about “weakly” and “strongly positive” forms or currents, we use the French convention about positivity: “positive” always means $\geqslant 0$. We follow [De, Ha].
Definition 2.5: Let $(V,I)$ be a $2n$-dimensional real vector space equipped with a complex structure operator $I$, $I^{2}=-\operatorname{Id}$. A weakly positive $(p,p)$-form on $V$ is a real $(p,p)$-form $\eta\in\Lambda^{p,p}_{\mathbb{R}}(V^{*})$ which satisfies
$\eta(x_{1},Ix_{1},x_{2},Ix_{2},\ldots,x_{p},Ix_{p})\geqslant 0$
for all $x_{1},\ldots,x_{p}\in V$.
Remark 2.6: Clearly, the set of weakly positive $(p,p)$-forms is a convex cone.
Definition 2.7: The cone of strongly positive polyvectors is the cone in $\Lambda^{p,p}_{\mathbb{R}}(V)$ generated by the monomials $x_{1}\wedge Ix_{1}\wedge x_{2}\wedge Ix_{2}\wedge\ldots\wedge x_{p}\wedge Ix_{p}$ with positive coefficients. Replacing $V$ with $V^{*}$, we also define weakly positive $(p,p)$-polyvectors and strongly positive $(p,p)$-forms.
Remark 2.8: By definition, the cone of strongly positive polyvectors is dual to the cone of weakly positive forms under the natural pairing $\Lambda^{p,p}_{\mathbb{R}}(V^{*})\times\Lambda^{p,p}_{\mathbb{R}}(V){\>\longrightarrow\>}{\mathbb{R}}$.
Remark 2.9: Fix a volume element $\operatorname{Vol}\in\Lambda^{2n}(V^{*})$ positive with respect to $I$. For any $v\in\Lambda^{m}(V)$, consider the contraction $v\,\lrcorner\,\operatorname{Vol}\in\Lambda^{2n-m}(V^{*})$. For any monomial $\eta\in\Lambda^{m}(V)$, $\eta\,\lrcorner\,\operatorname{Vol}$ is the complementary monomial. Therefore, the contraction takes the strongly positive cone in $\Lambda^{p,p}_{\mathbb{R}}(V)$ to the strongly positive cone in $\Lambda^{n-p,n-p}_{\mathbb{R}}(V^{*})$. Using this remark, it is not hard to prove the following basic claim.
Proposition 2.10:  Consider the natural non-degenerate pairing
$\Lambda^{p,p}(V^{*})\times\Lambda^{n-p,n-p}(V^{*})\to{\mathbb{R}},\quad(\alpha,\beta)\mapsto\frac{\alpha\wedge\beta}{\operatorname{Vol}}.$
Under this pairing the cone of strongly positive forms is dual to the cone of weakly positive forms. Moreover, if $\alpha\in\Lambda^{p,p}(V^{*})$ is weakly positive and $\beta\in\Lambda^{q,q}(V^{*})$ is strongly positive, then $\alpha\wedge\beta$ is weakly positive.
Proof: [De].
Lemma 2.11:  Let $\eta$ be a $(p,0)$-form. Then the form $(-1)^{\frac{p(p-1)}{2}}(\sqrt{-1}\>)^{p}\eta\wedge\overline{\eta}$ is weakly positive.
Proof: Clearly, a $(p,p)$-form $\alpha\in\Lambda^{p,p}(V^{*})$ is weakly positive iff its restriction to any $p$-dimensional $I$-invariant subspace $W\subset V$ is a non-negatively oriented volume form. However, $\eta|_{W}$ is a complex linear volume form, hence it can be written as $uz_{1}\wedge z_{2}\wedge\ldots\wedge z_{p}$, where $z_{i}\in\Lambda^{1,0}(W^{*})$ is a basis and $u\in\mathbb{C}$. Write $z_{i}:=x_{i}+\sqrt{-1}\>Ix_{i}$, where $x_{i}\in W^{*}$. Then $z_{i}\wedge\overline{z}_{i}=-2\sqrt{-1}\>x_{i}\wedge Ix_{i}$ which gives
$uz_{1}\wedge z_{2}\wedge...\wedge z_{p}\wedge\overline{u}\overline{z}_{1}\wedge\overline{z}_{2}\wedge...\wedge\overline{z}_{p}=(-1)^{\frac{p(p-1)}{2}}(-2\sqrt{-1}\>)^{p}|u|^{2}\prod_{i=1}^{p}x_{i}\wedge Ix_{i}.$
The volume form $\prod_{i=1}^{p}x_{i}\wedge Ix_{i}$ is positively oriented. This implies that $(-1)^{\frac{p(p-1)}{2}}(\sqrt{-1}\>)^{p}\eta\wedge\overline{\eta}$ is weakly positive.
2.4 Positive currents
Let $M$ be a compact complex manifold of complex dimension $n$. Denote by $\mathcal{E}^{k}(M)$ the space of complex $k$-forms of class $C^{\infty}$ on $M$. Fixing an arbitrary Hermitian metric and an affine connection on $M$, one may define the $C^{d}$-norm of $k$-forms, for arbitrary $d\geqslant 0$. It is well known that this family of norms makes the space of smooth $k$-forms $\mathcal{E}^{k}(M)$ a Fréchet topological vector space, the topology being independent on the Hermitian metric and the connection. The space $\mathcal{E}^{k}(M)$ has the Hodge decomposition
$\mathcal{E}^{k}(M)=\bigoplus_{p+q=k}\mathcal{E}^{p,q}(M),$
where $\mathcal{E}^{p,q}(M)$ are the closed subspaces of differential forms of type $(p,q)$. We will denote by $\mathcal{E}^{k}(M,\mathbb{R})$ the space of real $k$-forms with the induced topology.
Denote by $\mathcal{E}^{\prime}_{k}(M)$ the topological dual of $\mathcal{E}^{k}(M)$, i.e. the space of continuous linear forms on $\mathcal{E}^{k}(M)$. We will always work with the weak topology on $\mathcal{E}^{\prime}_{k}(M)$, i.e. the coarsest topology for which all the evaluation maps $\mathrm{ev}_{\alpha}\colon\mathcal{E}^{\prime}_{k}(M)\to\mathbb{C}$, $T\mapsto T(\alpha)$ for arbitrary $\alpha\in\mathcal{E}^{k}(M)$ are continuous. The topological vector space $\mathcal{E}^{\prime}_{k}(M)$ is called the space of currents of dimension $k$. The pairing $\mathcal{E}^{2n-k}(M)\times\mathcal{E}^{k}(M)\to\mathbb{C}$ given by the wedge product and integration over $M$ induces a natural continuous embedding $\mathcal{E}^{2n-k}(M)\hookrightarrow\mathcal{E}^{\prime}_{k}(M)$. For this reason the elements of $\mathcal{E}^{\prime}_{k}(M)$ are also called currents of degree $2n-k$.
The Hodge decomposition of $k$-forms induces the Hodge decomposition of currents. We denote by $\mathcal{E}^{\prime}_{p,q}(M)$ the dual of $\mathcal{E}^{p,q}(M)$ and call it the space of currents of bidimension $(p,q)$ or of bidegree $(n-p,n-q)$. The space of real currents will be denoted by $\mathcal{E}^{\prime}_{k}(M,\mathbb{R})$ and $\mathcal{E}^{\prime}_{p,p}(M,\mathbb{R})=\mathcal{E}^{\prime}_{p,p}(M)\cap\mathcal{E}^{\prime}_{2p}(M,\mathbb{R})$.
The de Rham differential $d\colon\mathcal{E}^{k}(M)\to\mathcal{E}^{k+1}(M)$ is continuous, and by the well known theory of elliptic complexes its image is closed in Fréchet topology. It follows from the Banach open mapping theorem and the Hahn–Banach theorem that the adjoint map $d\colon\mathcal{E}^{\prime}_{k+1}(M)\to\mathcal{E}^{\prime}_{k}(M)$ also has closed image, i.e. the space of exact currents is closed in the weak topology.
Recall that a real current $T\in\mathcal{E}^{\prime}_{1,1}(M,\mathbb{R})$ is called positive if for any $\eta\in\mathcal{E}^{1,0}(M)$ we have $\sqrt{-1}T(\eta\wedge\overline{\eta})\geqslant 0$. Let us denote the set of positive currents by $\mathcal{E}^{+}_{1,1}(M)$. This set is clearly a closed convex cone in $\mathcal{E}^{\prime}_{1,1}(M,\mathbb{R})$. The positive currents admit the following description, see e.g. [HL, §3]. Given a positive current $T\in\mathcal{E}^{+}_{1,1}(M)$ there exists a positive Borel measure $\|T\|$ called the total variation of $T$ and a $\|T\|$-measurable section $\tau$ of the bundle $\Lambda^{1,1}TM$, i.e. a $\|T\|$-measurable bivector field of Hodge type $(1,1)$, such that $T=\tau\|T\|$. The latter equality means that the action of $T$ on $\alpha\in\mathcal{E}^{1,1}(M)$ is given by
$T(\alpha)=\int_{M}\langle\tau,\alpha\rangle d\|T\|,$
where $\langle\tau,\alpha\rangle$ denotes the pairing between bivectors and two-forms. The bivector field $\tau$ is positive in the following sense: for any $\eta\in\mathcal{E}^{1,0}(M)$ the function $\langle\tau,\sqrt{-1}\eta\wedge\overline{\eta}\rangle$ is non-negative $\|T\|$-almost everywhere on $M$.
Proposition 2.12:  Let $\pi\colon M\to X$ be a proper holomorphic map, and $T\in\mathcal{E}^{+}_{1,1}(M)$ a positive current such that $T(\pi^{*}\alpha)=0$ for any positive $(1,1)$-form $\alpha$ on $X$. Then in the above decomposition $T=\tau\|T\|$ the bivector field $\tau$ is almost everywhere tangent to the fibers of $\pi$, i.e. $\tau\in\ker{d\pi}$ outside a subset of $\|T\|$-measure zero.
Proof: The differential $d\pi:\;T_{x}M{\>\longrightarrow\>}T_{\pi(x)}X$ can be interpreted as a map from $TM$ to $\pi^{*}(TX)$. Then $d\pi(\tau)$ is a measurable section of the bundle $\pi^{*}\Lambda^{1,1}TX$. By our assumptions, for any $\eta\in\mathcal{E}^{1,0}(X)$ we have
$\sqrt{-1}T(\pi^{*}\eta\wedge\pi^{*}\overline{\eta})=\int_{M}\langle d\pi(\tau),\sqrt{-1}\eta\wedge\overline{\eta}\rangle d\|T\|=0,$
and since $d\pi(\tau)$ is positive, the function $\langle d\pi(\tau),\sqrt{-1}\eta\wedge\overline{\eta}\rangle$ is non-negative almost everywhere on $M$, therefore it must vanish almost everywhere. Since this is true for arbitrary $\eta$ as above, we conclude that $d\pi(\tau)$ vanishes almost everywhere, which concludes the proof.
3 Hermitian symplectic structures
In the sequel, we use the following version of the Hahn–Banach theorem, sometimes called the “Hahn–Banach separation theorem”.
Theorem 3.1:  (Hahn–Banach theorem, [Bou], or [Sch, Chapter II, Theorem 3.1])
Let $V$ be a topological vector space, $A\subset V$ an open convex subset of $V$, and $W$ a subspace of $V$ satisfying $W\cap A=\emptyset$. Then there exists a continuous linear functional $\theta$ on $V$, such that $\theta|_{A}>0$ and $\theta|_{W}=0$.
Recall the following definitions.
Definition 3.2: Let $(M,I)$ be a complex manifold and $\omega\in\Lambda^{2}M$ a real symplectic form. The complex structure $I$ is tamed by $\omega$ if for any non-zero tangent vector $v$ at an arbitrary point of $M$ we have $\omega(v,Iv)>0$. In this case we call $\omega$ a Hermitian symplectic structure. Equivalently, a symplectic form $\omega$ is Hermitian symplectic if its $(1,1)$-part $\omega^{1,1}$ is a Hermitian form.
The following characterization of compact complex manifolds admitting Hermitian symplectic structures is originally due to Sullivan [Su], see [Ch], [DP] and [HL] for related results and a broader context.
Theorem 3.3:  Let $M$ be a compact complex manifold. The manifold $M$ does not admit a Hermitian symplectic structure if and only if if admits a non-zero positive exact $(n-1,n-1)$-current.
Proof: We sketch the well known proof for the sake of completeness and because we refer to these arguments elsewhere. We use the Hahn–Banach separation theorem (3), applied to $V$, $A$ and $W$ defined below.
Let $V=\Lambda^{2}(M,{\mathbb{R}})$ be the Fréchet space of smooth two-forms equipped with the $C^{\infty}$-topology, $A$ the cone of two-forms with strictly positive $(1,1)$-part, and $W\subset V$ the space of closed two-forms. The intersection $A\cap W$ is the set of all Hermitian symplectic forms. If this set is empty, we use 3 to find a non-zero functional $\theta\in V^{*}$ which is positive on $A$ (this means that $\theta$ is a positive $(n-1,n-1)$-current) and vanishing on $W$. The latter condition means that $\theta$ is exact. Indeed, $\langle\theta,du\rangle=0$ for all $u\in\Lambda^{1}M$ implies that $\theta$ is closed. It is exact, because otherwise it is non-zero on $W$ by Poincaré duality (the de Rham cohomology of currents is equal to the de Rham cohomology of differential forms, [De]).
Theorem 3.4:  Let $\pi\colon M\to B$ be a holomorphic Lagrangian fibration on a compact simply connected hyperkähler manifold of maximal holonomy, and $\mathcal{M}_{t}=(M,I_{t})$, $t\in{\mathbb{C}}$ the corresponding degenerate twistor deformation. Then $\mathcal{M}_{t}$ admits an Hermitian symplectic structure for all $t\in{\mathbb{C}}$.
Proof:  We use 3. Let $T\in\mathcal{E}^{+}_{1,1}(\mathcal{M}_{t})$ be an exact positive current of type $(1,1)$ on $\mathcal{M}_{t}$. We need to prove that $T=0$.
As explained above, $T$ admits a representation $T=\tau\mu$, where $\mu=\|T\|$ is a positive Borel measure and $\tau$ is a $\mu$-measurable bivector field on $\mathcal{M}_{t}$. Let us check that the conditions of 2.4 are satisfied. Fix a Kähler form $\eta\in\Lambda^{1,1}B$. For any positive form $\alpha\in\Lambda^{1,1}B$ we have $0\leqslant\alpha\leqslant C\eta$ for some constant $C>0$. Then $0\leqslant T(\pi^{*}\alpha)\leqslant T(\pi^{*}\eta)$, but $\pi^{*}\eta$ is closed, and since $T$ is exact $T(\pi^{*}\eta)=0$. Therefore $T(\pi^{*}\alpha)=0$ and we conclude by 2.4 that the bivector field $\tau$ is tangent to the fibers of $\pi$ (after, possibly, changing $\tau$ on a subset of $\mu$-measure zero).
Fix a point $x\in M$, let $V=T_{x}M\otimes\mathbb{C}$ and let $W=\ker(d\pi)_{x}\subset V$ be the complexified tangent space to the fiber of $\pi$ through $x$. The bivector $\tau(x)$ is a linear combination with positive coefficients of bivectors of the form $-\sqrt{-1}u\wedge\overline{u}$ for $u\in V^{1,0}_{t}$. But since $\tau$ is tangent to the fibers of $\pi$, we have $u\in W^{1,0}_{t}$ for all bivectors appearing in that linear combination. Now, we know by 2.2 that the subspace $W^{1,0}_{t}\subset W$ does not depend on $t$. We conclude that the bivector field $\tau$ is of type $(1,1)$ and positive for all complex structures $I_{t}$ in the degenerate twistor family. Therefore $T$ is an exact positive current of bidimension $(1,1)$ on $\mathcal{M}_{0}=M$. The latter being a Kähler manifold, the current $T$ must vanish by 3.
4 $d$-Commendable manifolds
[…]
Proposition 4.2:  Let $M$ be a complex manifold with a Hermitian symplectic structure $\omega$. Then $M$ is $d$-commendable, for all $d>0$.
Proof:  For $0<d<\dim_{\mathbb{C}}(M)$ let $\eta=\omega^{d}$. Then $\eta$ closed and $\eta^{d,d}$ is a positive linear combination of the forms $(\omega^{2,0}\wedge\overline{\omega^{2,0}})^{k}\wedge(\omega^{1,1})^{d-2k}$, for $k=0,\ldots,[d/2]$. The form $(\omega^{2,0})^{k}\wedge(\overline{\omega^{2,0}})^{k}$ is weakly positive by 2.3, and $(\omega^{1,1})^{d-2k}$ is strongly positive for $2k<d$. By 2.3 the forms $(\omega^{2,0}\wedge\overline{\omega^{2,0}})^{k}\wedge(\omega^{1,1})^{d-2k}$ are weakly positive. Moreover, $(\omega^{1,1})^{d}\geqslant Ch^{d}$ for some constant $C>0$, because $\omega^{1,1}$ is a Hermitian form. Therefore $\eta^{d,d}\geqslant Ch^{d}$.
4.1 Barlet spaces on commendable manifolds
The purpose of introducing the notion of $d$-commendability is that it can serve as a replacement of the condition of being Kähler in some situations. More precisely, we will observe that the Barlet spaces of $d$-dimensional cycles of $d$-commendable manifolds behave in some ways similar to the spaces of cycles of Kähler manifolds. Let us start with the following claim.
Proposition 4.3:  Let $M$ be a $d$-commendable compact complex manifold and $\mathcal{C}_{d}$ a connected component of the Barlet space of dimension $d$ cycles on $M$. Then $\mathcal{C}_{d}$ is compact.
Proof:  Let $h\in\Lambda^{1,1}M$ be a Hermitian metric on $M$ and $\eta$ a closed $2d$-form with $\eta^{d,d}\geqslant Ch^{d}$. By the compactness criterion of Barlet, based on Bishop’s theorem ([Mag], [BM, Theorem 4.2.69]), we need to show that the volume of all cycles in $\mathcal{C}_{d}$ with respect to $h$ is bounded. Assume that $[Z]\in\mathcal{C}_{d}$ is a cycle. Then
$\int_{Z}h^{d}\leqslant(1/C)\int_{Z}\eta^{d,d}=(1/C)\int_{Z}\eta.$
Since $d\eta=0$, the latter integral depends only on the cohomology class of $[Z]$ which is constant for all cycles in the connected component $\mathcal{C}_{d}$. The claim follows.
In what follows we will need a relative version of the above statement. We will consider the following setting. Let $\pi\colon\mathcal{M}\to B$ be a proper smooth holomorphic map between complex manifolds. Let $\mathcal{C}_{d}(\mathcal{M}/B)$ be a connected component of the Barlet space of $d$-dimensional cycles contained in the fibers of $\pi$, see [BM, section 4.8.2]. Then the map $p\colon\mathcal{C}_{d}(\mathcal{M}/B)\to B$ that sends a cycle to the point $t\in B$ over which this cycle is supported is holomorphic (loc. cit.) The following proposition is a generalization of [BM, Corollary 4.8.].
Proposition 4.4:  In the above setting assume that there exists a differential form $\eta\in\Lambda^{2d}(\mathcal{M})$ such that for any $t\in B$ the fiber $\mathcal{M}_{t}$ and the restriction $\eta_{t}=\eta|_{\mathcal{M}_{t}}$ satisfy the conditions of $d$-commendability. Then the support morphism $p\colon\mathcal{C}_{d}(\mathcal{M}/B)\to B$ is proper, in particular its image is an analytic subvariety of $B$.
Proof:  To prove properness we restrict to a compact subset $K\subset B$. The cycles in the preimage $p^{-1}(K)$ are supported in the compact subset $\pi^{-1}(K)$ of $\mathcal{M}$. So by the compactness criterion [BM, Theorem 4.2.69] we need to prove that the volume of the cycles $[Z]\in\mathcal{C}_{d}(\mathcal{M}/B)$ in $p^{-1}(K)$ is bounded above. Let us fix a Hermitian metric $h$ on $\mathcal{M}$.
We can assume that $K$ is sufficiently small and find a constant $C>0$ such that for any $t\in K$ we have $\eta_{t}^{d,d}\geqslant Ch_{t}^{d}$, where the subscript $t$ denotes the restriction to the fiber $\mathcal{M}_{t}$. Now, given a cycle $[Z]\in p^{-1}(K)$ let $t\in K$ be the point over which $[Z]$ is supported. We estimate the volume of $[Z]$:
$\int_{Z}h^{d}=\int_{Z}h^{d}_{t}\leqslant(1/C)\int_{Z}\eta_{t}^{d,d}=(1/C)\int_{Z}\eta_{t}.$
The homology class of $[Z]$ is constant, and the restriction of $\eta$ to any fiber $\mathcal{M}_{t}$ is closed, so the integral $\int_{Z}\eta_{t}$ depends only on the point $t\in K$ over which $[Z]$ is supported; let us denote the value of the integral by $\varphi(t)$. Since $\eta$ is a smooth differential form on $\mathcal{M}$, $\varphi$ is a continuous function. Since $K$ is compact, $\varphi$ is bounded from above on $K$. This gives the required upper bound for the volume of the cycles $[Z]\in p^{-1}(K)$.
4.2 Non-separated points in the Teichmüller space
Let $M$ be a simply connected compact hyperkähler manifold of maximal holonomy and $\pi\colon M\to B$ a holomorphic Lagrangian fibration. Let $\mathcal{M}\to\mathbb{A}^{1}$ be the corresponding degenerate twistor family. Denote by $U\subset\mathbb{A}^{1}$ the open subset of points of the base such that for $t\in U$ the fiber $\mathcal{M}_{t}$ is Kähler.
The argument used to prove 4.2 below is a version of the one used in [Hu1] to establish the bimeromorphic correspondence between non-separated fibers of a twistor deformation. In [Hu1], both fibers are assumed to be Kähler. We prove the same statement when the fibers are Hermitian symplectic.
By the global Torelli theorem ([V1]) the hyperkähler manifold $\mathcal{M}_{t}$ admits a hyperkähler deformation $X$ such that the period of $X$ corresponds to the point $t_{0}$ and $X$ lies in the same connected component of the Teichmüller space as $\mathcal{M}_{t}$ ([V3, Section 1.2]). Let $\mathcal{X}\to\Delta$ denote the restriction of the universal deformation of $X$ to $\Delta$.
Proposition 4.5:  Consider a family $\mathcal{M}\to\mathbb{A}^{1}$ of holomorphically symplectic manifolds, and let $U\subset\mathbb{A}^{1}$ be the set of all points corresponding to Kähler fibers. Assume the the fiber over a limit point $t_{0}\in\partial U$ is Hermitian symplectic. Then there exists a hyperkähler deformation $X$ of $M$ that is bimeromorphic to $\mathcal{M}_{t_{0}}$
Proof:  Let $\Delta\subset\mathbb{A}^{1}$ be a small disc around $t_{0}$. Note that $\Delta\cap U\neq\emptyset$, so we may find a Kähler fiber $\mathcal{M}_{t}$ with $t\in\Delta$.
Let $\omega$ be a Hermitian symplectic form on $\mathcal{M}_{t_{0}}$, and $\omega^{\prime}$ a Kähler form on $X$.
Note that $\omega+\omega^{\prime}$ is a Hermitian symplectic form on $X\times\mathcal{M}_{t_{0}}$. By continuity, after possibly shrinking $\Delta$, we may assume that the restriction of $\omega+\omega^{\prime}$ to any fiber $\mathcal{X}_{t}\times\mathcal{M}_{t}$, $t\in\Delta$ makes this fiber Hermitian symplectic.
The idea now is to take the graph $\Gamma_{t}$ of the bimeromorphism $\mathcal{X}_{t}\dashrightarrow\mathcal{M}_{t}$ and use 4.1 to show that there is a subsequence $\{\Gamma_{t_{i}}\}$ converging to a subvariety $\Gamma_{t_{0}}$ in $\mathcal{X}_{t_{0}}\times\mathcal{M}_{t_{0}}$. Then we use a topological argument to show that one of the irreducible components of $\Gamma_{t_{0}}$ is a graph of a bimeromorphism $\mathcal{X}_{t_{0}}\dashrightarrow\mathcal{M}_{t_{0}}$.
Let us consider the open subset $\Delta\cap U$. It is non-empty because $t_{0}$ is a limit point of $U$. The fibers of the family $\mathcal{M}$ over this open subset are Kähler by the definition of $U$. The points of the Teichmüller space corresponding to $\mathcal{M}_{t}$ and $\mathcal{X}_{t}$, $t\in\Delta\cap U$ are either equal or non-separated by construction, hence by the theorem of Huybrechts [Hu1, Theorem 4.3] the manifolds $\mathcal{M}_{t}$ and $\mathcal{X}_{t}$ are bimeromorphic for any $t\in\Delta\cap U$.
Consider the complex manifold $\mathcal{Y}=\mathcal{X}\times_{\Delta}\mathcal{M}$. For $t\in\Delta\cap U$ the graph of a bimeromorphic map between $\mathcal{X}_{t}$ and $\mathcal{M}_{t}$ is a subvariety $\Gamma_{t}\subset\mathcal{Y}$. Let $\mathcal{C}_{\dim_{\mathbb{C}}X}(\mathcal{Y}/\Delta)$ be an irreducible component of the relative Barlet space that contains a point corresponding to one of these graphs. By 4.1 the image of $\mathcal{C}_{\dim_{\mathbb{C}}X}(\mathcal{Y}/\Delta)$ in $\Delta$ is a closed subvariety, hence it is either a countable set of points, or the whole disc $\Delta$. Since the subset $\Delta\cap U$ is uncountable, one of the connected components of the Barlet space must map surjectively onto $\Delta$.
We conclude that there exists a sequence of cycles $[\Gamma_{t_{i}}]$ in the fibers $\mathcal{Y}_{t_{i}}$, where $t_{i}\in\Delta\cap U$, $t_{i}\to t_{0}$ for $i\to\infty$, giving the above bimeromorphic maps, contained in the same irreducible component of the Barlet space and converging to a limit cycle $[\Gamma_{t_{0}}]$ in the fiber $X\times\mathcal{M}_{t_{0}}$ over $t_{0}$. Then the homology classes of $[\Gamma_{t_{i}}]$ do not depend on $i$. Denote by $\Omega^{\prime}$ a holomorphic symplectic form on $X$. The bimeromorphic maps induced by $\Gamma_{t_{i}}$ map a holomorphic symplectic form on $\mathcal{M}_{t_{i}}$ to a holomorphic symplectic form on $\mathcal{X}_{t_{i}}$. Therefore by continuity $[\Gamma_{t_{0}}]$ seen as a correspondence acting on cohomology maps the cohomology class $[\Omega_{t_{0}}]$ on $\mathcal{M}_{t_{0}}$ to the cohomology class $[\Omega^{\prime}]$ on $X$. We can therefore apply 4.2 below to conclude that the cycle $[\Gamma_{t_{0}}]$ contains exactly one irreducible component that is the graph of a bimeromorphic map between $\mathcal{M}_{t_{0}}$ and $X$.
The proof of the following proposition is a version of [Hu1, proof of Theorem 4.3]. We use it to establish the bimeromorphic equivalence of two holomorphically symplectic manifolds, the only difference from loc. cit. being that we do not assume one of the manifolds to be Kähler.
We consider the following setting. Assume that $M$ is a compact complex manifold with a holomorphic symplectic form $\Omega$ and $X$ is a compact hyperkähler manifold of maximal holonomy with a holomorphic symplectic form $\Omega^{\prime}$. Assume that $\dim_{\mathbb{C}}(M)=\dim_{\mathbb{C}}(X)=2n$ and that $[Z]$ is a $2n$-cycle on $X\times M$. Assume that $[Z]$ is of degree one both over $X$ and over $M$. Assume moreover that $[Z]_{*}\colon H^{2}(M,\mathbb{C})\to H^{2}(X,\mathbb{C})$ is an isomorphism that sends $[\Omega]$ to $[\Omega^{\prime}]$. We can decompose the cycle into a finite sum $[Z]=\sum a_{i}[Z_{i}]$, where $a_{i}>0$ and $Z_{i}\subset X\times M$ is an irreducible subvariety.
Proposition 4.6:  In the above decomposition of $[Z]$ there is exactly one irreducible component dominating both $M$ and $X$ and this component is the graph of a bimeromorphic isomorphism between $M$ and $X$.
Proof:  The degree of $[Z]=\sum a_{i}[Z_{i}]$ over $X$ is equal to $\sum a_{i}\mathrm{deg}_{X}(Z_{i})$, and since all $a_{i}$ are positive, there exists $i_{0}$ such that $\mathrm{deg}_{X}(Z_{i_{0}})=1$, $a_{i_{0}}=1$ and $\mathrm{deg}_{X}(Z_{i})=0$ for $i\neq i_{0}$. Analogously, there exists $i_{1}$, such that $\mathrm{deg}_{M}(Z_{i_{1}})=1$, $a_{i_{1}}=1$ and $\mathrm{deg}_{M}(Z_{i})=0$ for $i\neq i_{1}$. We need to prove that $i_{0}=i_{1}$: in this case the subvariety $Z_{i_{0}}=Z_{i_{1}}$ is the graph of a bimeromorphic map between $X$ and $M$.
Denote by $p\colon X\times M\to X$ and $q\colon X\times M\to M$ the two projections. Let us consider the action of the correspondence given by $[Z]$ on the cohomology class of the symplectic form $\Omega$. Recall that by our assumptions $[Z]_{*}[\Omega]=[\Omega^{\prime}]$. For a connected component $Z_{i}$ of the cycle $Z$ denote by $\widetilde{Z}_{i}$ its resolution of singularities and by $\widetilde{p}_{i}\colon\widetilde{Z}_{i}\to X$ and $\widetilde{q}_{i}\colon\widetilde{Z}_{i}\to M$ the induced maps. Then $[Z_{i}]_{*}[\Omega]$ is the cohomology clases of the current $\widetilde{p}_{i*}(\widetilde{q}_{i}^{*}\Omega)$. For $i\neq i_{0}$ the map $\widetilde{p}_{i}$ is not dominant, and since $\widetilde{q}_{i}^{*}\Omega$ is a holomorphic two-form, the current $\widetilde{p}_{i*}(\widetilde{q}_{i}^{*}\Omega)$ vanishes. Indeed, a closed (2,0)-current is holomorphic, hence smooth by elliptic regularity, but the support of $\widetilde{p}_{i*}(\widetilde{q}_{i}^{*}\Omega)$ is contained in a proper subvariety. Now consider the case $i=i_{0}$. Then $\widetilde{p}_{i_{0}*}(\widetilde{q}_{i_{0}}^{*}\Omega)$ is a holomorphic two-form on $X$, therefore it must be proportional to $\Omega^{\prime}$, because $X$ is hyperkähler of maximal holonomy. It follows that $\widetilde{q}_{i_{0}}^{*}\Omega$ is generically a symplectic form, so $\widetilde{q}_{i_{0}}$ must be dominant, hence $i_{0}=i_{1}$.
4.3 Degenerate twistor deformations are Kähler
We finish the paper by proving our main result.
Theorem 4.7:  Let $M$ be a simply connected compact hyperkähler manifold of maximal holonomy and $\pi\colon M\to B$ a holomorphic Lagrangian fibration. Let $\mathcal{M}\to\mathbb{A}^{1}$ be the corresponding degenerate twistor family. Then all fibers $\mathcal{M}_{t}$, $t\in\mathbb{A}^{1}$ are Kähler.
Proof:  Denote by $U\subset\mathbb{A}^{1}$ the open subset of points of the base such that for $t\in U$ the fiber $\mathcal{M}_{t}$ is Kähler. Assume that $U\neq\mathbb{A}^{1}$ and let $x\in\partial U$. Then by 4.2, the fiber $\mathcal{M}_{t}$ is of Fujiki class C. Moreover, by 3 the manifold $\mathcal{M}_{t}$ admits a Hermitian symplectic form $\omega$. Then $\omega^{1,1}$ is a Hermitian metric and $dd^{c}(\omega^{1,1})=0$, therefore $\mathcal{M}_{t}$ is Kähler by [Ch, Theorem 0.2]. [^2] We thank the participants of the mathoverflow discussion for this reference It follows that $U=\mathbb{A}^{1}$.
Acknowledgements: We are grateful to Ekaterina Amerik, Alexey Golota, Evgeny Shinder, Olivier Martin and Eyal Markman for interesting discussions of the subject.